OpenStudy (anonymous):

Find the slope of the line whose equation is 17 + 3y = 7x

OpenStudy (anonymous):

First you have to change it to slope form (y=mx+b).

OpenStudy (the_fizicx99):

You want to change 17 + 3y = 7x into y = mx + b So first start by subtracting 17 from both sides

OpenStudy (anonymous):

3y=-17+7x?

OpenStudy (anonymous):

Yes but we need into slope form (y=mx+b) so lets make it 3y=7x-17. Now you want to divide both sides by 3.

OpenStudy (anonymous):

y=7x-17/3?

OpenStudy (anonymous):

So is it now 17/3?

OpenStudy (anonymous):

You divide 17/3 and 7x/3 @tHe_FiZiCx99

OpenStudy (the_fizicx99):

Why did you tag me so state you were going to divide? ._.

OpenStudy (anonymous):

So is the slope 17/3 or 7/3?

OpenStudy (anonymous):

the slope would be 2.33 and the y-intercept is 5.66

OpenStudy (anonymous):

17+3y=7x Since 17 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 17 from both sides. 3y=-17+7x Move all terms not containing y to the right-hand side of the equation. 3y=7x-17 Divide each term in the equation by 3. (3y)/(3)=(7x)/(3)-(17)/(3) Simplify the left-hand side of the equation by canceling the common factors. y=(7x)/(3)-(17)/(3) Combine the numerators of all expressions that have common denominators. y=(7x-17)/(3)

OpenStudy (anonymous):

my options are 17/3 7/3 -7/3

OpenStudy (anonymous):

it would then be 7/3